PREALGEBRA: Multiplication and Division of Mixed Numerals; Applications
There are about 155,000 Labrador retrievers registered with The American Kennel Club. This is 3 and 4/9 times the number of beagles registered. How many beagles are registered? Answer by schrammbledeggs(41) 
 Ok so there are 155,000 Labradors and the number of Beagles is 3 and 4/9 times that amount...to turn a mixed number into a fraction take the denominator and multiply it by the whole number...then add the numerator to that number and stick that in the numerator and keep your original denominator...
So...9 (denominator) * 3 (whole number) = 27
27 + 4 (numerator) = 31 (new numerator)
Keep old denominator...new fraction is...
31/9
Now since 150,000 is 31/9 times the number of beagles...we must divide 150,000 by 31/9...
But instead of doing division...we can find out the same thing by multiplying by the reciprocal of 31/9...which is 9/31
so 150,000 * 9/31 = 43,548 beagles (approximately)
:)